# Flavino Dark Matter in a Non-Abelian Discrete Flavor Model

### v1

@@
Takaaki Nomura $1$nomura@scu.edu.cn,  Yusuke Shimizu 2,3222shimizu.yusuke@kaishi-pu.ac.jp,  Towa Takahashi 2333towa@muse.sc.niigata-u.ac.jp
@@
We study a relic density of the flavino dark matter in modified Altarelli and Feruglio $A_{4}$ model which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. We calculate the Lagrangian from the superpotential in the model. In estimating relic density, we consider the relevant interactions from the Lagrangian that realize the vacuum expectation value alignments and charged lepton masses where we assume that the supersymmetry breaking effects are small for flavon sector. As a result, we find the degenerate masses between the lightest “flavon” and “flavino”, and only two parameters in the potential determines the relic density. Then allowed parameter space of these parameters are estimated from relic density calculation taking a constraint from lepton flavor violation into account. We also briefly discuss other dark matter physics such as direct detection, indirect detection and collider search. )
@@ 1 Introduction
The flavor symmetry can be applied to the three generations of the SM fermions. The Froggatt-Nielsen mechanism [3] which was proposed by C. D. Froggatt and H. B. Nielsen applies a global $U(1)_{FN}$ symmetry. Thanks to the symmetry, one can naturally explain the fermion mass hierarchies. The non-Abelian discrete symmetry (See for the review [4]-[8].) can naturally explain the lepton mixing angles which are tri-bimaximal mixing (TBM) [9, 10] before the reactor experiments reported the non-zero reactor angle $\theta_{13}$ [1, 2]. After measurements of the $\theta_{13}$, many authors have studied the breaking or deviation from the TBM [11]-[33] or other patterns of the lepton mixing angles, e.g. tri-bimaximal-Cabibbo mixing [34, 35]. One of the successful flavor models was proposed by G. Altarelli and F. Feruglio [36, 37]. The Altarelli and Feruglio (AF) model realizes the TBM by using the non-Abelian discrete symmetry $A_{4}$. They mainly introduced two SM gauge singlet scalar fields so-called “flavons” taking the vacuum expectation value (VEV) alignments of the $A_{4}$ triplets as $(1,0,0)$ and $(1,1,1)$, which naturally explains that the charged lepton mass matrix is diagonal and neutrino mixing is TBM, respectively. However, the correct VEV alignments cannot be driven from the potential analysis. Then, they adopted the supersymmetry (SUSY) and introduced the so-called “driving” fields.
@@
In this paper, we discuss a relic density of the flavino DM in modified AF $A_{4}$ model [27],[48]-[51] which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. In estimating relic density, we consider relevant interactions from superpotential that realize the VEV alignments and charged lepton masses where we assume SUSY breaking effects are small for flavon sector. As a result, we find degenerate mass between the lightest flavon and flavino, and only two parameters in the potential determines the relic density. Then allowed parameter space of these parameters are estimated from relic density calculation taking a constraint from lepton flavor violation (LFV) into account. We also briefly discuss other DM physics such as direct detection, indirect detection and collider search.
This paper is organized as follows. In Section 2, we briefly introduce the modified AF $A_{4}$ model which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. In Section 3, we discuss the relic density of the flavino DM calculating its annihilation cross section. Section 4 is devoted to a summary. We show the basis and relevant multiplication rule of the $A_{4}$ group in Appendix A. The LFV constraint for our model is shown in Appendix B.
@@ 2 Model
In this section, we briefly introduce an $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ model [27],[48]-[51]. Three generations of the left-handed lepton doublet superfields $\Phi_{\ell}=(\Phi_{\ell 1},\Phi_{\ell 2},\Phi_{\ell 3})$ are assigned to the triplet as $\bf 3$ for $A_{4}$ symmetry. The right-handed charged lepton superfields $\Phi_{e_{R}^{c}}$, $\Phi_{\mu_{R}^{c}}$, and $\Phi_{\tau_{R}^{c}}$ are assigned to the singlets as $\bf 1$, $\bf 1^{\prime\prime}$, and $\bf 1^{\prime}$ for $A_{4}$ symmetry, respectively. One of the Higgs doublet superfields is defined as $\Phi_{d}$ and assigned to the singlet as $\bf 1$ for $A_{4}$ symmetry. We introduce a superfields $\Phi_{T}=(\Phi_{T1},\Phi_{T2},\Phi_{T3})$ which are so-called “flavon superfields” and assigend to the triplet as $\bf 3$ for $A_{4}$ symmetry. In order to obtain the relevant couplings, we add the $Z_{3}$ symmetry. We also add the $U(1)_{R}$ symmetry so that we can generate the VEVs and VEV alignments through the $F$-terms by coupling flavon superfields to so-called “driving superfields” $\Phi_{0}^{T}=(\Phi_{01}^{T},\Phi_{02}^{T},\Phi_{03}^{T})$ which are assigned to triplet as $\bf 3$ for $A_{4}$ symmetry and carry the $R$ charge $+2$ under $U(1)_{R}$ symmmetry. The charge assignments of $SU(2)_{L}$, $A_{4}$, $Z_{3}$, and $U(1)_{R}$ are summarized in Tab. 1.
@@
where "$c"$ indicates charge conjugation and $X$ is identified as Dirac fermion. Here we focus on interactions among the DM $X$ and the lightest flavons $\varphi_{T1}$ and $\phi^{T}_{01}$ since it provide dominant contributions in relic density estimation as we discuss in next section. The relevant interactions are summarized by
@@
These interactions induce DM annihilation processes that determine relic density.
@@ 3 Dark matter physics
In this section we discuss relic density of flavino DM calculating its annihilation cross section. Other DM phenomenologies such as direct/indirect detections and collider searches are also discussed briefly.
@@ 3.1 DM annihilation cross sections
In our scenario flavino mainly annihilate into the lightest flavons, $\{\varphi_{T},\phi_{01}^{T}\}$, due to small couplings among flavino and the SM leptons. There are many possible processes of such annihilation processes as

### v2

@@
Takaaki Nomura $1$nomura@scu.edu.cn,  Yusuke Shimizu 2,3222shimizu.yusuke@kaishi-pu.ac.jp, and  Towa Takahashi 2333towa@muse.sc.niigata-u.ac.jp
@@
We study a relic density of the “flavino” dark matter in modified Altarelli and Feruglio $A_{4}$ model which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. We calculate the Lagrangian from the superpotential in the model. In estimating the relic density, we consider the relevant interactions from the Lagrangian that realize the vacuum expectation value alignments and charged lepton masses where we assume that the supersymmetry breaking effects are small for “flavon” sector. As a result, we find the degenerate masses between the lightest “flavon” and “flavino”, and only two parameters in the potential determines the relic density. Then the allowed parameter space of these parameters are estimated from the relic density calculation taking a constraint from lepton flavor violation into account. We also briefly discuss other dark matter physics such as the direct detection, indirect detection, and collider search. )
@@ 1 Introduction
The flavor symmetry can be applied to the three generations of the SM fermions. The Froggatt-Nielsen mechanism [3] which was proposed by C. D. Froggatt and H. B. Nielsen applies a global $U(1)_{FN}$ symmetry. Thanks to the symmetry, one can naturally explain the fermion mass hierarchies. The non-Abelian discrete symmetry (See for the review [4]-[8].) can naturally explain the lepton mixing angles which are tri-bimaximal mixing (TBM) [9, 10] before the reactor experiments reported the non-zero reactor angle $\theta_{13}$ [1, 2]. After measurements of the $\theta_{13}$, many authors have studied the breaking or deviation from the TBM [11]-[33] or other patterns of the lepton mixing angles, e.g. tri-bimaximal-Cabibbo mixing [34, 35]. One of the successful flavor models was proposed by G. Altarelli and F. Feruglio [36, 37]. The Altarelli and Feruglio (AF) model realizes the TBM by using the non-Abelian discrete symmetry $A_{4}$. They mainly introduced two SM gauge singlet scalar fields so-called “flavons”. Taking the vacuum expectation value (VEV) alignments of the $A_{4}$ triplets as $(1,0,0)$ and $(1,1,1)$, one can naturally explain that the charged lepton mass matrix is diagonal and neutrino mixing is TBM, respectively. However, the correct VEV alignments cannot be driven from the potential analysis. Then, they adopted the supersymmetry (SUSY) and introduced the so-called “driving” fields [36, 37].
@@
In this paper, we discuss a relic density of the flavino DM in modified AF $A_{4}$ model [27],[48]-[51] which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. In estimating the relic density, we consider relevant interactions from superpotential that realize the VEV alignments and charged lepton masses where we assume SUSY breaking effects are small for flavon sector. As a result, we find degenerate mass between the lightest flavon and flavino, and only two parameters in the potential determines the relic density. Then allowed parameter space of these parameters are estimated from relic density calculation taking a constraint from lepton flavor violation (LFV) into account. We also briefly discuss other DM physics such as the direct detection, indirect detection and collider search.
This paper is organized as follows. In Section 2, we briefly introduce the modified AF $A_{4}$ model which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. In Section 3, we discuss the relic density of the flavino DM calculating its annihilation cross section. Section 4 is devoted to summary. We show the relevant multiplication rule of the $A_{4}$ group in Appendix A. The LFV constraint for our model is shown in Appendix B.
@@ 2 Model
In this section, we briefly introduce an $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ model [27],[48]-[51]. Three generations of the left-handed lepton doublet superfields $\Phi_{\ell}=(\Phi_{\ell 1},\Phi_{\ell 2},\Phi_{\ell 3})$ are assigned to the triplet as $\bf 3$ for $A_{4}$ symmetry. The right-handed charged lepton superfields $\Phi_{e_{R}^{c}}$, $\Phi_{\mu_{R}^{c}}$, and $\Phi_{\tau_{R}^{c}}$ are assigned to the singlets as $\bf 1$, $\bf 1^{\prime\prime}$, and $\bf 1^{\prime}$ for $A_{4}$ symmetry, respectively. One of the Higgs doublet superfields is defined as $\Phi_{d}$ and assigned to the singlet as $\bf 1$ for $A_{4}$ symmetry. We introduce superfields $\Phi_{T}=(\Phi_{T1},\Phi_{T2},\Phi_{T3})$ which are so-called “flavon superfields” and assigend to the triplet as $\bf 3$ for $A_{4}$ symmetry. In order to obtain the relevant couplings, we add the $Z_{3}$ symmetry. We also add the $U(1)_{R}$ symmetry so that we can generate the VEVs and VEV alignments through the $F$-terms by coupling flavon superfields to so-called “driving superfields” $\Phi_{0}^{T}=(\Phi_{01}^{T},\Phi_{02}^{T},\Phi_{03}^{T})$ which are assigned to triplet as $\bf 3$ for $A_{4}$ symmetry and carry the $R$ charge $+2$ under $U(1)_{R}$ symmmetry. The charge assignments of $SU(2)_{L}$, $A_{4}$, $Z_{3}$, and $U(1)_{R}$ are summarized in Tab. 1.
@@
where "$c"$ indicates charge conjugation and $X$ is identified as Dirac fermion. Here we focus on interactions among the DM $X$ and the lightest flavons $\varphi_{T1}$ and $\phi^{T}_{01}$ since it provide dominant contributions in the relic density estimation as we discuss in next section. The relevant interactions are summarized by
@@
These interactions induce DM annihilation processes that determine the relic density.
@@ 3 Dark matter physics
In this section we discuss the relic density of flavino DM calculating its annihilation cross section. Other DM phenomenologies such as direct/indirect detections and collider searches are also discussed briefly.
@@ 3.1 DM annihilation cross sections
In our scenario the flavinos mainly annihilate into the lightest flavons, $\{\varphi_{T},\phi_{01}^{T}\}$, due to small couplings among flavino and the SM leptons. There are many possible processes of such annihilation processes as